OpenStudy (anonymous):

(x^4+y^5)(x^4-y^5)... same as my last question, I can work it out with numbers but just with variables throws me off....please help!!!!

OpenStudy (anonymous):

Multiplying?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you familiar with the factored form of a 'difference of squares?"

OpenStudy (anonymous):

no not at all

OpenStudy (anonymous):

ok, let's look at something simpler then.. What would you get if you multiplied \((a+b)(a-b)?\)

OpenStudy (anonymous):

a^2-b^2 ??

OpenStudy (anonymous):

Right, so you see you have a difference of squares there: one square (a^2) minus another square (b^2). In your above example, you have something very similar.

OpenStudy (anonymous):

So do the same as with (a+b)(a-b), but now a = x^4 and b = y^5.

OpenStudy (anonymous):

but what about the exponents how do I work those

OpenStudy (anonymous):

When you multiplied a*a = a^2, it was really a^1 * a^1 = a^2. If you multiply powers of the same base, you add the exponents, so for x^4 * x^4, it is x^(4+4). Understand?

OpenStudy (anonymous):

I think...ok... so I wouldn't add the two 4's???